# TinyLLM: Learning a Small Student from Multiple Large Language Models

### v2

@@ 4.2 Performance Comparison
Table 1: Overall experimental results. The best results across different datasets and LLM sizes are highlighted in bold. $\Delta_{FF}$ and $\Delta_{Distill}$ represent the relative performance improvement of TinyLLM to Full Fine-Tuning and Distill-step-by-step, respectively. Accuracy is used as the evaluation metric.
Comparison to Baselines Methods. The results of six datasets and two reasoning tasks are shown in Table  1. From the table, we observe that the employment of a full fine-tuning method, despite its theoretically enhanced capacity for parameter adjustment, does not consistently yield superior results to LoRA. Conversely, TinyLLM demonstrates substantial performance enhancements across all datasets and LLM sizes. Quantitatively, TinyLLM secures an average performance improvement by +15.69%, +11.55%, and +5.07% compared to full fine-tuning for 80M, 250M, and 780M student models, respectively. Compared to state-of-the-art distillation method Distill-step-by-step, TinyLLM achieves an improvement of +10.00%, +10.32%, and +11.79% correspondingly. This demonstrates the effectiveness of TinyLLM and highlights the advantages of the designs in our method.
Comparison to Teachers. TinyLLM also shows superior performance compared to the teacher models. For example, a 780M student can achieve an average performance of 73.88 across different datasets, which is +14.56% better than the 3B teacher and +23.40% better than the 7B teacher. Moreover, an even smaller student model with 250M parameters can outperform the teachers (+0.82% to 3B, +8.60% to 7B) while using only 8.3% and 3.6% of the teacher parameters.
4.3 Efficiency Analysis of Training Set Size in Knowledge Transfer
Advantage Over Standard Knowledge Distillation. With the aim of fully evaluating the proposed model, we test the performance of TinyLLM against state-of-the-art across different sizes of training sets. As shown in Figure 2, TinyLLM demonstrates superior performance in comparison to the state-of-the-art Distill-step-by-step method when using various ratios of the training data. In addition, we observe that when using more training data, the performance improves. Notably, in certain scenarios, our model, even when trained with just 12.5% of the training set size, can surpass the performance of Distill-step-by-step, which uses a considerably larger dataset. This is evident in the case of the PQA dataset, where TinyLLM achieves performance that is comparable to or even exceeds that of Distill-step-by-step, despite being trained on a much smaller portion of the dataset. This observation holds true for both the 80M and 250M student models, indicating the superiority of our model across models of different sizes.
Outperforming full fine-tuning. TinyLLM further demonstrates its superiority compared to full fine-tuning using the whole dataset. Remarkably, in the situations of training a 250M TinyLLM on ARC and PQA datasets, alongside training an 80M TinyLLM on the PQA dataset, only 12.5% of the training data is required to exceed the benchmarks set by full fine-tuning. In addition, when training an 80M TinyLLM on the ARC dataset, a 75% reduction in training samples is adequate for TinyLLM to attain a higher accuracy than what is achieved through full fine-tuning on the whole dataset.
@@ 4.3 Efficiency Analysis of Training Set Size in Knowledge Transfer
Figure 2: A comparative analysis of TinyLLM against the state-of-the-art Distill-step-by-step method using 80M and 250M FLAN-T5 model architectures across various training set sizes. The dotted line indicates the full fine-tuning ($F$) using 100% dataset. It is evident that TinyLLM consistently surpasses the performance of both Distill-step-by-step and full fine-tuning. Notably, TinyLLM achieves this superior accuracy while employing substantially fewer training examples.
4.4 Ablation Study
For a comprehensive evaluation, we conduct ablation studies to validate the contribution of the in-context example generator and rationales from multiple teachers in enhancing the reasoning capabilities of the distilled LLM. To facilitate this, we create three ablation variants of TinyLLM, each designed to assess the impact of specific components on the overall performance:
  • w/o in-context rules out the use of in-context examples during rationale generation. This variant tests the efficacy of in-context examples in guiding the student model to generate more accurate and relevant rationales.
  • w/o LLaMa and w/o T5 exclude the rationale supervision from the corresponding teacher model during distillation. These variants help in understanding the individual contributions of each teacher’s rationales in enriching the student model’s learning experience and overall performance.
The results, as detailed in Table 2, provide a comparative analysis between the ablation variants and the complete TinyLLM model. From Table 2, we have the following observations: (1) TinyLLM outperforms all ablation variants. (2) There is no significant performance gap between the ablation variants. These observations confirm the hypothesis that absorbing high-quality rationales from multiple teacher models significantly contributes to the refinement of the reasoning capabilities of the distilled LLM. The observations also imply a balanced importance between the in-context examples and the supervision of multiple teachers.
@@ 4.4 Ablation Study
Table 2: Impact of in-context examples and contributions of various teachers.
Figure 3: Performance w.r.t. different values of weight $\alpha$.
4.5 Parameter Sensitivity
To fully evaluate the model, we conduct parameter sensitivity experiments in Figure  3 for commonsense reasoning dataset ARC and biomedical reasoning dataset PQA. In particular, we focus on the exploration of trade-off weights $\alpha^{T5}$ and $\alpha^{LLaMA}$, which reveals the model’s adaptability across different choices of parameter values. Based on the figure, we can derive the following observations: 1) the optimal parameters for various datasets and tasks differ. The reason for this phenomenon is that biomedical reasoning questions are often lengthy and complex, weakening the impact of rationales from teachers and making a small value of $\alpha$ sufficient. In contrast, commonsense reasoning questions are typically concise and straightforward, making the rationales from teacher models valuable and leading to a large value of $\alpha$. 2) Increasing $\alpha$ generally improves performance, attributed to the Chain-of-Thought reasoning of the student model. While accuracy may not directly reflect reasoning quality, the prediction capability benefits from the multi-task learning process. However, excessively high $\alpha$ values degrade the performance by shifting the focus from prediction to reasoning. 3) The sensitivity of $\alpha^{T5}$ and $\alpha^{LLaMA}$ varies across datasets. For commonsense reasoning, $\alpha^{T5}$ is more sensitive than $\alpha^{LLaMA}$, whereas, for biomedical reasoning, their sensitivities are similar, indicating that rationales from various teacher models can have different insights and contributions to the particular task.
4.6 Case Study
To analyze the underlying reason for the superiority of TinyLLM in a more intuitive perspective, we perform case studies by comparing the predictions of different models. Table  3 presents two examples we randomly selected from the ARC and PIQA datasets.
In the first example, T5 provides a completely incorrect rationale while LLaMA generates a meaningful rationale. Accordingly, T5 has a wrong prediction while LLaMA predicts correctly. Full fine-tuning results in an incorrect answer and the state-of-the-art Distill-step-by-step method also predicts incorrectly. The reason behind this is that the distillation method, which utilizes T5’s reasoning, introduces noises that misguide the student model, even though T5 significantly outperforms LLaMA in accuracy on the ARC dataset (as shown in Table 1). However, TinyLLM demonstrates its ability to infer the correct answer (a) with the guidance of rationales from both teachers.
In the second example, LLaMA, Distill-step-by-step, and full fine-tuning all produce the incorrect prediction. Although LLaMA achieves a significantly better performance on the PIQA dataset compared to the T5 (i.e., accuracy 78.80 compared to 58.43 according to Table 1), LLaMA still returns an incorrect rationale. This indicates that the reliance on a single-teacher model can be misleading. On the other hand, TinyLLM can generate the correct prediction (a). These examples underscore the significance of multi-teacher learning, as proposed in our TinyLLM framework, which provides the student model with diverse contextual insights essential for effective learning and reasoning.
@@ 4.6 Case Study
Table 3: Case study of different models’ prediction. Examples are selected from the ARC and PIQA datasets. In both cases, TinyLLM successfully generates the correct answer.
5 Conclusion
@@
In this paper, we propose TinyLLM, a novel knowledge distillation paradigm to learn a small student LLM from multiple large teacher LLMs. TinyLLM involves several principled designs, such as learning contextually appropriate rationales using an in-context example generator, enabling the credibility of rationales with a teacher-forcing Chain-of-Thought strategy, and inheriting a wider range of knowledge from various teachers. Our extensive empirical evaluation and in-depth analysis, conducted across six datasets spanning two reasoning tasks, demonstrate that TinyLLM brings significant and consistent improvements by up to $15.69\%$ over full fine-tuning, up to +23.40% over teacher models, and up to +11.79% over state-of-the-art. Moreover, TinyLLM holds a significantly smaller model size, e.g., 1.1% to 26.0% compared to the sizes of the teachers.

### v3

@@ 3.3. Learning a Small Student
4. Experiments
In this section, we rigorously test TinyLLM against a series of empirical benchmarks across varied datasets and reasoning tasks. In addition, we conduct efficiency analyses, ablation studies, parameter sensitivity tests, and case studies to demonstrate the effectiveness and superiority of our method.
4.1. Experimental Setup
Datasets. For the task of commonsense reasoning, we use OpenBookQA (OBQA) (Mihaylov et al., 2018), The AI2 Reasoning Challenge (ARC) (Clark et al., 2018), Physical Interaction Question Answering (PIQA) (Bisk et al., 2020), and RiddleSense (Riddle) (Lin et al., 2021). For the task of biomedical reasoning, we consider PubMedQA (PQA) (Jin et al., 2019) and BioASQ (Tsatsaronis et al., 2015).
Baselines. We benchmark TinyLLM against the teacher’s performance and various baseline methods, including Inference-only that only leverage the pre-trained model for evaluation without training, and multiple fine-tuning methods that provide further adaptation. In particular, we consider LoRA (Hu et al., 2022), full fine-tuning, and the PINTO method (Wang et al., 2022a) for the fine-tuning methods. We also compare TinyLLM with various knowledge distillation strategies. To illustrate, we include standard KD (Hinton et al., 2015) that enforces the student to mimic the teacher’s labels and the Distill-step-by-step method (Hsieh et al., 2023) that leverage rationales.
Implementation Details. For all distillation baselines and TinyLLM, we set the learning rate to $5\times 10^{-5}$, batch size to 8, maximum input length to 1024, and epoch to 1. For Distill-step-by-step and TinyLLM, the trade-off weights $\alpha_{T_{n}}$ are explored within {0.01, 0.1, 0.5, 1, 2, 3}. We report the best result for Distill-step-by-step by leveraging different teacher models. For the choice of LLMs, we use FLAN-T5 (Chung et al., 2022) small (80M), base (250M), and large (780M) as the student, and FLAN-T5 xlarge (3B) and LLaMA 2-chat (Touvron et al., 2023b) (7B) as teachers. Experiments are conducted on four NVIDIA H100 Tensor Core GPUs.
4.2. Performance Comparison
Comparison to Baselines Methods. We conducted a thorough evaluation of our method, TinyLLM, across six diverse datasets spanning two distinct reasoning tasks: commonsense reasoning and biomedical reasoning. The detailed results are presented in Table  1, offering a comprehensive view of the performance landscape across different model sizes and methodologies. From the table, we derive several observations as follows.
First, while full fine-tuning is theoretically capable of maximizing parameter adjustments and should yield the best results in principle, it does not consistently outperform parameter-efficient fine-tuning methods such as LoRA. This outcome suggests that simply having a larger number of adjustable parameters does not guarantee improved generalization or performance, especially when the fine-tuning process may inadvertently overfit the training data or fail to capture nuanced task-specific knowledge.
Second, TinyLLM demonstrates consistent and substantial performance improvements across all datasets and model sizes, underscoring the robustness and adaptability of our approach. Specifically, the quantitative gains achieved by TinyLLM are noteworthy: we observe an average performance boost of +15.69%, +11.55%, and +5.07% for student models with 80M, 250M, and 780M parameters, respectively, compared to full fine-tuning. These improvements are significant, particularly considering that larger models often exhibit diminishing returns on performance gains, making these results even more compelling.
Third, when compared to the state-of-the-art distillation method, Distill-step-by-step, TinyLLM achieves impressive relative improvements of +10.00%, +10.32%, and +11.79% for the 80M, 250M, and 780M student models, respectively. These results highlight the effectiveness of our method in distilling knowledge from large models into the smaller student model. The consistent superiority of TinyLLM across different model sizes and datasets demonstrates the advantages of our design choices, including the strategic balance of integrating intermediate rationale guidance and mimicking teacher outputs. Furthermore, using rationales during training enhances the model’s reasoning capability and improves interpretability over conventional distillation.
@@ 4.2. Performance Comparison
Figure 2. A comparative analysis of TinyLLM against the state-of-the-art Distill-step-by-step method using 80M and 250M FLAN-T5 model architectures across various training set sizes. Dotted line indicates the full fine-tuning (FF) using 100% dataset. It is evident that TinyLLM consistently surpasses the performance of both Distill-step-by-step and full fine-tuning. Notably, TinyLLM achieves this superior accuracy while employing substantially fewer training examples.
Comparison to Teachers. In addition to outperforming baseline methods, TinyLLM also exhibits superior performance compared to the original teacher models. This is particularly remarkable given the significant difference in model size. For instance, a 780M parameter student model trained with TinyLLM achieves an impressive average performance score of 73.88 across various datasets. This score represents a substantial improvement of +14.56% over the performance of 3B teacher model, and an even more striking +23.40% performance improvement over the 7B teacher model. However, further reducing the size of the student model can introduce difficulties for the student to learn and result in decreased performance.
These results demonstrate the efficacy of TinyLLM in transferring knowledge and suggest that our approach enables smaller models to generalize better and perform tasks more effectively than their larger counterparts. Furthermore, the efficiency gains are particularly pronounced when considering smaller student models, such as the 250M parameter model, which manages to surpass both the 3B and 7B parameter teachers with relative improvements of +0.82% and +8.60%, respectively. It is also worth mentioning that the 250M model operates with only 8.3% and 3.6% of the teacher models’ parameters, showcasing the remarkable efficiency of TinyLLM in compressing and enhancing the performance of smaller models without compromising accuracy.
4.3. Efficiency Analysis of Training Set Size in Knowledge Transfer
Advantage Over Standard Knowledge Distillation. To thoroughly assess the efficiency and effectiveness of our proposed method, TinyLLM, we conducted a series of experiments that evaluate its performance across varying training set sizes, particularly in comparison to the state-of-the-art Distill-step-by-step method. This analysis is crucial for understanding how well our model performs not only with full datasets but also under conditions of limited training data, which is often a real-world constraint.
As illustrated in Figure  2, TinyLLM consistently outperforms the Distill-step-by-step method across all tested ratios of the training data, demonstrating its robustness and efficiency. This superior performance is particularly evident even when we reduce the size of the training set. The trend indicates that TinyLLM can effectively leverage smaller amounts of data to achieve comparable or better results than methods that require larger training datasets.
A striking example of this efficiency is observed in the context of the PQA dataset. Here, TinyLLM, when trained with only 12.5% of the available training data, not only meets but often exceeds the performance levels achieved by Distill-step-by-step when utilizing the full training set. This significant reduction in required training data, without sacrificing performance, underscores the practical advantages of TinyLLM, particularly in scenarios where data is scarce or expensive to obtain.
This finding holds across different model sizes, including both 80M and 250M parameter student models. The consistent outstanding performance of TinyLLM in these cases suggests that our method is highly effective in transferring knowledge efficiently, making it a versatile tool for various applications, regardless of the model size. The capability of achieving superior performance with limited data can be particularly beneficial in resource-constrained environments, where computational resources and training time are restricted.
Outperforming Full Fine-Tuning. Beyond its advantages over standard knowledge distillation methods, TinyLLM also demonstrates significant improvements compared to traditional full fine-tuning approaches, even when using the entire dataset. This comparison further highlights the efficiency of our method in terms of both data usage and computational resources.
In particular, when training a 250M parameter model with TinyLLM on the ARC and PQA datasets, as well as training an 80M parameter model on the PQA dataset, only 12.5% of the full training data is necessary to surpass the benchmarks established by full fine-tuning. This result is remarkable, as it shows that TinyLLM can achieve superior performance with a fraction of the data typically required for full fine-tuning, significantly reducing the computational cost and time required for training.
In addition, for 80M student model trained on the ARC dataset, TinyLLM can achieve a higher accuracy than full fine-tuning with a 75% reduction in training samples, while further increasing the reduction to 87.5% results in a competitive performance. This result demonstrates the capability of TinyLLM in reducing the need for extensive training data to maintain or further improve the model performance, making it an attractive alternative to full fine-tuning.
Table 2. Impact of in-context examples, the teacher-forcing strategy and contributions of various teachers.
@@ 4.3. Efficiency Analysis of Training Set Size in Knowledge Transfer
4.4. Ablation Study
To provide a comprehensive evaluation of our proposed method, TinyLLM, we conducted an ablation study in Table 2 to validate the contributions of key components in enhancing the reasoning capabilities of the distilled LLM. Specifically, we focused on assessing the impact of the in-context example generator, the use of rationales from multiple teacher models, and the teacher-forcing strategy. By isolating these components, we aim to understand their individual and collective contributions to the overall performance of TinyLLM. We designed four ablation variants of TinyLLM, each purposefully modified to test the significance of specific components:
  • w/o in-context removes the use of in-context examples during rationale generation. This variant is crucial for evaluating the role of in-context examples in guiding the student model to produce more accurate, relevant, and contextually appropriate rationales.
  • w/o LLaMA and w/o T5 exclude the rationale supervision provided by the respective teacher models during the distillation process. By removing the influence of either LLaMA or T5, these variants help us understand the individual contributions of each teacher’s rationales.
  • w/o diverse teachers excludes the weaker teacher model, instead generating multiple rationales from the stronger teacher model. This variant is designed to test the effectiveness of using a diverse set of teachers, as opposed to relying on a single, potentially more robust teacher.
  • w/o teacher-forcing eliminates the teacher-forcing strategy during rationale generation. Teacher-forcing is a technique where the model is trained using the correct output (from the teacher) as input for the next step, rather than its own previous prediction. By removing this strategy, we aim to assess its effectiveness in helping the student model generate higher-quality rationales.
Table 2 provides a comparative analysis between each ablation variant and the complete TinyLLM model. From the data presented in Table 2, several important observations emerge: (1) TinyLLM consistently outperforms all ablation variants, demonstrating that the integration of all components—rationales from multiple teachers, in-context examples, and the teacher-forcing strategy—collectively contributes to the model’s superior performance. This result highlights the synergistic effect of these components, where their combination leads to significant improvements in the reasoning capabilities of the distilled LLM. (2) There is no substantial performance gap between the different ablation variants, suggesting that while each component contributes to the model’s performance, none of them alone is solely responsible for the observed improvements. This finding implies a balanced importance among the components, with each playing a complementary role in refining the model’s reasoning abilities.
4.5. Parameter Sensitivity
To thoroughly evaluate the robustness and adaptability of our proposed model, we conducted parameter sensitivity experiments on two datasets in different tasks: ARC for commonsense reasoning and PQA for biomedical reasoning. These experiments are critical for understanding how different parameter settings, particularly the trade-off weights $\alpha^{T5}$ and $\alpha^{LLaMA}$, influence the model’s performance across various tasks. The results of these sensitivity analyses are depicted in Figure  3.
In this analysis, our primary focus is on exploring the effects of varying the trade-off weights $\alpha^{T5}$ and $\alpha^{LLaMA}$, which balance the influence of rationales provided by the T5 and LLaMA teacher models, respectively. This exploration reveals the model’s adaptability and how it responds to different parameter configurations, offering insights into the optimal settings for maximizing performance across different tasks and datasets. From the results shown in Figure  3, our key observations are as follows.
Optimal Parameter Variability Across Datasets and Tasks. It is evident that the optimal parameter settings for $\alpha^{T5}$ and $\alpha^{LLaMA}$ vary depending on the dataset and the nature of the reasoning task. For biomedical reasoning tasks, such as those found in the PQA dataset, questions tend to be lengthy and complex, often requiring deep comprehension and retrieval of detailed information. In these cases, the impact of rationales from the teacher models is partially diminished, given the complexity of the content overshadows the direct utility of rationale guidance. As a result, a smaller value of $\alpha$ is sufficient to achieve optimal performance. On the other hand, commonsense reasoning tasks, exemplified by the ARC dataset, typically involve more concise and straightforward questions. In these scenarios, the rationales provided by the teacher models play a more critical role in guiding the student model’s reasoning process, leading to the need for a larger $\alpha$ value to fully leverage this guidance.
Effect of Increasing $\alpha$ Values on Performance. Increasing the values of $\alpha^{T5}$ and $\alpha^{LLaMA}$ generally leads to improved performance, which can be attributed to the enhanced Chain-of-Thought reasoning capabilities of the student model. By placing greater emphasis on the rationales from teacher models, the student model benefits from a richer multi-task learning experience, which enhances its prediction capabilities. However, this improvement has its limits. Excessively high values of $\alpha$ can degrade performance by causing the model to overly focus on reasoning processes at the expense of prediction accuracy. This shift in focus may lead to a situation where the model becomes too reliant on rationales, potentially overfitting to the rationale structure rather than learning to generalize effectively from the data.
Differential Sensitivity of $\alpha^{T5}$ and $\alpha^{LLaMA}$ Across Datasets. The sensitivity of the parameters $\alpha^{T5}$ and $\alpha^{LLaMA}$ varies between different datasets, reflecting the diverse contributions of the T5 and LLaMA teacher models to the reasoning process. For commonsense reasoning tasks, we observe that $\alpha^{T5}$ is more sensitive than $\alpha^{LLaMA}$. This suggests that the rationales generated by the T5 model are particularly valuable for tasks involving straightforward reasoning, where the logical flow and concise reasoning provided by T5 are more impactful. Conversely, in biomedical reasoning tasks, the sensitivities of $\alpha^{T5}$ and $\alpha^{LLaMA}$ are more balanced, indicating that both teacher models offer valuable, albeit different, insights that complement each other in processing the complex and detailed content typical of biomedical texts. This balance highlights the importance of leveraging diverse teacher models to capture the full range of reasoning required for different tasks.
Figure 3. Performance w.r.t. different values of weight $\alpha$.
4.6. Case Study
@@
Figure 4. Case study of different models’ prediction. Examples are selected from the ARC and PIQA datasets. In both cases, TinyLLM successfully generates the correct answer.
To gain a more intuitive understanding of why TinyLLM consistently outperforms other models, we conducted case studies by comparing the predictions generated by different models. These case studies offer valuable insights into the specific scenarios where TinyLLM excels and highlight the advantages of our multi-teacher approach. Figure  4 presents two representative examples that we randomly selected from the ARC and PIQA datasets, illustrating the differences in model predictions and the underlying rationales.
In the first example, taken from the ARC dataset, we observe a significant discrepancy in the reasoning capabilities of the teacher models. The T5 model provides a completely incorrect rationale, leading to an incorrect prediction. In contrast, LLaMA generates a meaningful and accurate rationale, which correctly guides the prediction. However, the full fine-tuning method, which does not utilize teacher rationales, also fails to provide the correct answer. Similarly, the state-of-the-art Distill-step-by-step method also predicts incorrectly, likely because the distillation process, which leverages T5’s reasoning, introduces noise that misguides the student model. This occurs even though T5 significantly outperforms LLaMA in overall accuracy on the ARC dataset (as shown in Table 1). However, TinyLLM demonstrates its robustness by correctly inferring the answer (a), effectively understanding the rationales from both T5 and LLaMA, and filtering out the noise introduced by the incorrect rationale from T5. This example highlights the strength of TinyLLM in balancing and integrating multiple sources of knowledge, leading to superior prediction accuracy.
A similar situation occurs in the second example from the PIQA dataset. T5 provides an incorrect rationale, while LLaMA aligns with the ground truth. Other methods failed to predict the correct answer. However, TinyLLM successfully integrates and balances rationales from both teachers.
5. Conclusion and Future Work
@@
In this paper, we propose TinyLLM, a novel knowledge distillation paradigm to learn a small student LLM from multiple large teacher LLMs. TinyLLM involves several principled designs, such as learning contextually appropriate rationales using an in-context example generator, enabling the credibility of rationales with a teacher-forcing Chain-of-Thought strategy, and inheriting a wider range of knowledge from various teachers. Our extensive empirical evaluation and in-depth analysis, conducted across six datasets spanning two reasoning tasks, demonstrate that TinyLLM brings significant and consistent improvements by up to +15.69% over full fine-tuning, up to +23.40% over teacher models, and up to +11.79% over state-of-the-art. Moreover, TinyLLM holds a significantly smaller model size, e.g., 1.1% to 26.0% compared to the sizes of the teachers.
As future work, we envision several promising directions to further enhance TinyLLM. One potential avenue is to manage any conflicting rationales that can arise when integrating a broader set of teacher LLMs. Conflicting rationales can emerge due to hallucinations, variations in reasoning styles, or domains of expertise for different teacher models. Another direction involves using open-source embeddings to represent questions within each dataset. This approach maps questions into a semantic space and allows us to select in-context examples based on similarity or the top nearest neighbors. By incorporating targeted in-context example selection, we can assess whether providing such focused and meaningful context can enhance the comprehension and performance of the student model.